prove (cosx^3-sinx^3)/(cosx-sinx)=(cosx^2)/(sinx+cosx)
Answer by Alan3354(69443)

prove (cosx^3-sinx^3)/(cosx-sinx)=(cosx^2)/(sinx+cosx)
---------
I'll assume you mean cos^3(x), etc, not cos(x^3)
---
(cosx^3-sinx^3)/(cosx-sinx)=(cosx^2)/(sinx+cosx)
cos^2 + sin*cos + sin^2 = cos^2/(sin+cos)
1 + sin*cos = cos^2/(sin+cos)
=====================
At x = 45 degs:
1 + 1/2 =? (1/2)/sqrt(2)
Not an identity.